OpenStudy (anonymous):

18/45. reduce to lowest terms.

OpenStudy (ashleyisakitty):

divide by three and find that your answer is 2/5

OpenStudy (zzr0ck3r):

divide top and bottom by 9

OpenStudy (anonymous):

Thanks Ashley.:)

OpenStudy (whpalmer4):

write each number as its factors: \[\frac{18}{45} = \frac{2*3*3}{3*3*5} = \frac{2*\cancel{3*3}}{\cancel{3*3}*5} = \frac{2}{5}\]